Wow. Everyone must be on vacation this week. This'll be the 3rd time I'm posting this one, so far no answer. help, it's due by tomorrow... =)

Ride the peaks. Smith bicycled 45 miles going east from Durango, and Jones bicycled 70 miles. Jones averaged 5 miles per hour more than Smith, and his trip took one-half hour longer than Smith’s. How fast was each one traveling?

Thanks

Answer by checkley71(8403)

45/X+.5=70/(X+5) OR (45+.5X)/X=70/(X+5) OR 70X=45X+.5X~2+225+2.5X OR
.5X~2-22.5X+225=O OR X~2-45X+450)OR (X-15)(X-30)=0 OR X=15 & X=30
PROOF 45/15+.5=70/(15+5) OR 3+.5=3.5 OR 3.5=3.5
SMITHS SPEED WAS 15 MPH & JONES WAS 15+5=20 MPH.